OpenStudy (anonymous):

how to find the sum of the whole numbers from 1 to 100. with Carl Friedrich Gauss`es method

OpenStudy (anonymous):

1+100=101 2+99=101 3+98=101 ... Gauss realized that you get the sum of 101, 50 times. Thus the sum of the whole numbers 1 to 100 is 101 x 50 = 5050